# Decoupling Knowledge and Reasoning in Transformers: A Modular Architecture with Generalized Cross-Attention

### v1

@@ Abstract
Transformers have achieved remarkable success across diverse domains, but their monolithic architecture presents challenges in interpretability, adaptability, and scalability. This paper introduces a novel modular Transformer architecture that explicitly decouples knowledge and reasoning through a generalized cross-attention mechanism to a shared knowledge base, specifically designed for effective knowledge retrieval. Critically, we provide a rigorous mathematical derivation demonstrating that the Feed-Forward Network (FFN) in a standard Transformer is a specialized case (a closure) of this generalized cross-attention, revealing its role in implicit knowledge retrieval and validating our design. This theoretical framework provides a new lens for understanding FFNs and lays the foundation for future research exploring enhanced interpretability, adaptability, and scalability, enabling richer interplay with external knowledge bases and other systems.
@@ 1 Introduction
Large language models (LLMs) based on the Transformer architecture have achieved remarkable success in various natural language processing tasks [19, 6, 14, 3]. However, their monolithic architecture, where knowledge and reasoning are deeply intertwined, presents critical challenges for real-world applications requiring transparency, adaptability, and continuous learning.
@@
  • Interpretability. Understanding how Transformers arrive at conclusions is difficult [4, 20]. The distributed nature of knowledge representation, particularly within Feed-Forward Networks (FFNs), hinders pinpointing the information used for specific predictions, limiting their use in high-stakes applications like healthcare and legal reasoning.
  • Adaptability. Adapting pre-trained models to new knowledge or integrating them with external systems is inefficient and complex. Current methods like RAG [13], which simply concatenate retrieved context with input, suffer from context dilution and static, input-level retrieval, often resulting in insufficient or excessive information. This hinders continuous learning and the ability to incorporate rapidly changing information, such as real-time news or scientific discoveries.
  • Scalability. The tight coupling of knowledge and reasoning in monolithic Transformers prevents independent scaling, leading to disproportionate parameter growth and hindering large-scale and continuous learning, unlike the additive scaling observed in human cognition, where people can efficiently transfer existing reasoning capabilities to new knowledge. This "parameter explosion" limits accessibility of larger models, hindering progress towards truly comprehensive knowledge-driven AI.
To address these limitations, we propose a novel modular Transformer architecture that explicitly decouples knowledge and reasoning through a generalized cross-attention mechanism to a shared knowledge base. Our core goal is to enable seamless interaction with external knowledge bases, facilitating continuous learning, knowledge sharing, and independent scaling of knowledge capacity.
A key contribution of this work is a rigorous theoretical analysis demonstrating that the FFN in a standard Transformer can be expressed as a specialized case (a closure) of our generalized cross-attention. This reveals the implicit knowledge retrieval role of FFNs and provides a crucial validation of our proposed mechanism. By establishing this functional equivalence under joint training (where the knowledge base is trained within the model), we provide a solid theoretical foundation for future exploration of external knowledge base integration. Because of this proven equivalence, we expect identical performance in this joint training setting. This theoretical groundwork is essential for the subsequent exploration of external knowledge integration, which is the primary focus of future work.
@@
  • A Novel Modular Architecture. We propose a modular Transformer architecture that explicitly decouples knowledge and reasoning through generalized cross-attention to a shared knowledge base.
@@ 3 A Modular Architecture with Explicit Knowledge Decoupling
To address these limitations, we propose a novel modular architecture (illustrated in Figure 1 (b)) that explicitly decouples reasoning and knowledge by introducing a shared knowledge base $E\in\mathbb{R}^{|E|\times d_{E}}$ across all layers, where $|E|$ is the number of knowledge entries and $d_{E}$ is the dimensionality of each entry. $E$ is accessed via dedicated cross-attention mechanisms. It is important to note that, in this work, we focus on the theoretical framework and the case where $E$ is trained jointly within the model. The exploration of external, pluggable knowledge bases is left for future work.
@@ 3.1 The Proposed Architecture
Our modular Transformer block replaces the standard Feed-Forward Network (FFN) with a cross-attention layer that attends to $E$. This design choice is motivated by the hypothesis that FFNs in standard Transformers implicitly perform a form of context-dependent knowledge retrieval. By making this implicit process explicit, we aim to gain significant advantages in terms of interpretability, adaptability, and scalability in future work with external knowledge bases.
Similar to the multi-head self-attention layer in a standard Transformer, each attention head within our modular blocks learns distinct projection matrices ($W_{Q}^{l}$, $W_{K}^{l}$, $W_{V}^{l}$), effectively focusing on different aspects or subsets of information within $E$. This mechanism enables dynamic, in-context knowledge retrieval, allowing the model to access and integrate relevant information from $E$ at each layer of processing.
@@ 7 Related Work
Transformer Architectures. The Transformer architecture [19] has revolutionized NLP and other fields, leading to various modifications and extensions for improved efficiency and performance [18, 15] and influential architectures like BERT [6] and GPT [14]. While much work has focused on the attention mechanism, the role of Feed-Forward Networks (FFNs) has received less attention. Geva et al. [8] addressed this by proposing that FFNs function as key-value memories, implicitly storing and retrieving knowledge. Our work builds directly on this insight, providing a formal mathematical derivation demonstrating that FFNs are a specialized case of generalized cross-attention. This formalization, further supported by empirical analyses of FFNs in code language models by Haider et al. [10], offers a deeper understanding of the functional role of FFNs.
Knowledge Retrieval and Symbolic AI. Integrating external knowledge into neural networks is a long-standing goal in bridging symbolic and neural AI [7]. Approaches like Memory Networks [21] and Neural Turing Machines [9] introduced explicit memory components accessed through attention. Retrieval-augmented language models (RAG) directly incorporate external knowledge into the input context. In contrast to these explicit methods, and building on the understanding of FFNs as implicit knowledge stores [8], our work proposes an architecture that explicitly decouples knowledge and reasoning, bridging the gap between implicit and explicit knowledge representation.
Modular Neural Networks. Modularity in neural networks has been shown to improve learning, generalization, and interpretability [12]. Previous work has explored task-specific modularity [1, 11] and parameterized Transformers [22], introducing modularity at a higher level (e.g., different modules for different tasks). Our work focuses on modularity within the Transformer architecture, formalizing the FFN as a module dedicated to implicit knowledge retrieval via cross-attention. This formalization lays the groundwork for explicitly decoupling knowledge into a separate module, distinguishing our approach from previous modular neural network designs.
Interpretability of Neural Networks. Various techniques, such as attention visualization [2], saliency maps [17], and probing tasks [5], aim to improve the interpretability of neural networks. While these methods provide insights into input importance, our work offers a theoretical framework for understanding the internal computations of FFNs, revealing their role as key-value memories [8]. This understanding, in conjunction with empirical analyses like those of Haider et al. [10], can inform more targeted interpretability methods, such as analyzing the folded weights in our derived formulation.

### v2

@@ Abstract
Transformers have achieved remarkable success across diverse domains, but their monolithic architecture presents challenges in interpretability, adaptability, and scalability. This paper introduces a novel modular Transformer architecture that explicitly decouples knowledge and reasoning through a generalized cross-attention mechanism to a globally shared knowledge base with layer-specific transformations, specifically designed for effective knowledge retrieval. Critically, we provide a rigorous mathematical derivation demonstrating that the Feed-Forward Network (FFN) in a standard Transformer is a specialized case (a closure) of this generalized cross-attention, revealing its role in implicit knowledge retrieval and validating our design. This theoretical framework provides a new lens for understanding FFNs and lays the foundation for future research exploring enhanced interpretability, adaptability, and scalability, enabling richer interplay with external knowledge bases and other systems.
@@ 1 Introduction
Large language models (LLMs) based on the Transformer architecture have achieved remarkable success [17, 7, 14, 3]. However, a key limitation of their monolithic design is the deep entanglement of knowledge and reasoning within the model’s parameters, posing significant challenges for real-world applications requiring transparency, adaptability, and continuous learning.
@@
  • Interpretability. The distributed nature of knowledge representation, especially within Feed-Forward Networks (FFNs), makes it difficult to pinpoint the specific information used for predictions [5, 18], limiting their use in high-stakes applications like healthcare and legal reasoning.
  • Adaptability. Adapting pre-trained models to new knowledge or integrating them with external systems is inefficient and complex. Current methods like RAG [12], which simply concatenate retrieved context with input, suffer from context dilution and retrieval performed only at the input level, lacking dynamic, context-dependent access to knowledge during processing. This hinders continuous learning and the ability to incorporate rapidly changing information, such as real-time news or scientific discoveries.
  • Scalability. The tight coupling of knowledge and reasoning in monolithic Transformers prevents modular scaling, unlike human cognition, which efficiently transfers existing reasoning capabilities to new knowledge through modular learning. This "parameter explosion" limits accessibility of larger models, hindering progress towards truly comprehensive knowledge-driven AI.
To address these limitations, while the interpretation of FFNs as implicit key-value memories [8] has provided valuable insights, a new challenge emerges when we seek to explicitly decouple knowledge and reasoning: how to maintain global knowledge consistency while enabling contextualized access at each layer. Current approaches, including those derived from the key-value perspective, often struggle to reconcile these two competing demands. To address this gap, we propose a novel modular Transformer architecture that explicitly decouples knowledge and reasoning through a generalized cross-attention mechanism to a globally shared knowledge base with layer-specific transformations. Our core goal is to enable seamless interaction with external knowledge bases, facilitating continuous learning, knowledge sharing, and independent scaling of knowledge capacity.
A key contribution of this work is a rigorous theoretical analysis demonstrating that the FFN in a standard Transformer can be expressed as a specialized case (a closure) of our generalized cross-attention. This reveals the implicit knowledge retrieval role of FFNs and provides a crucial validation of our proposed mechanism. By establishing this functional equivalence under joint training (where the knowledge base is trained within the model), we provide a solid theoretical foundation for future exploration of external knowledge base integration. Due to this proven equivalence, we expect functionally identical performance in this joint training setting. This theoretical groundwork is essential for the subsequent exploration of external knowledge integration, which is the primary focus of future work.
@@
  • A Novel Modular Architecture. We propose a modular Transformer architecture that explicitly decouples knowledge and reasoning through generalized cross-attention to a shared knowledge base with layer-specifc transformations.
@@ 3 A Modular Architecture with Explicit Knowledge Decoupling
To address the limitations of monolithic Transformers, particularly the entanglement of knowledge and reasoning (as discussed in Section 2), we propose a novel modular architecture. This architecture is designed to explicitly decouple knowledge and reasoning by introducing a dedicated mechanism for knowledge access.
3.1 Requirements for Explicit Knowledge Base
To effectively integrate knowledge into Transformer architectures, we believe a valid knowledge base (KB) must satisfy two key requirements:
  • Global sharing. A globally shared KB ensures parameter efficiency by avoiding redundant storage of the same knowledge across multiple layers. If each layer had its own KB, the number of parameters would increase significantly, leading to larger and more computationally expensive models. Second, a shared KB promotes knowledge consistency by ensuring that all parts of the model are working with the same information, preventing contradictions and inconsistencies. Finally, a centralized KB facilitates efficient management of knowledge, as updating, maintaining, and managing a single KB is much simpler than managing multiple distributed KBs, which is crucial for continuous learning and incorporating new information.
  • Layer-specific views. Different layers of a Transformer process information at different levels of abstraction. Therefore, each layer needs to access a specific view of the KB that is relevant to its current context. Analogous to how different areas of the human brain process information in distinct ways, requiring different perspectives on the same underlying knowledge, different layers of a Transformer require different perspectives on the shared KB.
@@ 3.2 The Proposed Architecture
To satisfy the requirements outlined above, our modular architecture introduces a globally shared knowledge base $E\in\mathbb{R}^{|E|\times d_{E}}$ across all layers with layer-specific transformations, where $|E|$ is the number of knowledge entries and $d_{E}$ is the dimensionality of each entry. $E$ is accessed via dedicated cross-attention mechanisms (illustrated in Figure 1 (b)). This effectively decomposes the implicit knowledge encoded within FFNs, resulting significant advantages in terms of interpretability, adaptability, and scalability in future work with external knowledge bases. It is important to note that, in this work, we focus on the theoretical framework and the case where $E$ is trained jointly within the model. The exploration of external, pluggable knowledge bases is left for future work.
Specifically, similar to the multi-head self-attention layer in a standard Transformer, each attention head within our modular blocks learns distinct projection matrices ($W_{Q}^{l}$, $W_{K}^{l}$, $W_{V}^{l}$), effectively focusing on different aspects or subsets of information within $E$. This mechanism enables dynamic, in-context knowledge retrieval, allowing the model to access and integrate relevant information from $E$ at each layer of processing.
@@ 7 Related Work
Transformers, FFNs, and Knowledge. A central challenge in integrating knowledge into Transformers is satisfying the two key requirements for a valid knowledge base (KB) discussed in Section 3.1: global sharing and layer-specific views. Existing approaches struggle to reconcile these competing demands. Geva et al. [8] proposed that FFNs function as implicit key-value memories, but this knowledge is inherently layer-local, violating the global sharing requirement and leading to redundancy and inconsistencies. Several subsequent approaches attempt to address scaling and knowledge management, but ultimately fail to simultaneously satisfy both requirements for a valid KB. PlugLM [4] introduces a shared knowledge base, satisfying the global sharing requirement, but uses identical keys and values for all layers, thus failing to provide layer-specific views. TokenFormer [13] tokenizes layer-specific weights as a form of compressed knowledge, again violating the global sharing requirement.
In contrast, our work directly addresses this tension by introducing a globally shared knowledge base with layer-specific transformations, achieved through a novel generalized cross-attention mechanism specifically designed for knowledge retrieval. This design uniquely satisfies both key requirements: a single, explicitly accessible knowledge base is shared across all layers (addressing parameter efficiency, knowledge consistency, and efficient management), while layer-specific transformations, implemented via distinct projection matrices and knowledge-dependent biases within our generalized cross-attention mechanism, ensure each layer accesses a unique, contextually relevant view of the knowledge (enabling more nuanced knowledge utilization).
Modular Neural Networks. Modularity in neural networks has been shown to improve learning, generalization, and interpretability [11]. Previous work has explored task-specific modularity [1, 10] and parameterized Transformers [19], introducing modularity at a higher level (e.g., different modules for different tasks). Our work focuses on modularity within the Transformer architecture, formalizing the FFN as a module dedicated to implicit knowledge retrieval via cross-attention. This formalization lays the groundwork for explicitly decoupling knowledge into a separate module (the globally shared KB), distinguishing our approach from previous modular neural network designs.
Interpretability of Neural Networks. Various techniques, such as attention visualization [2], saliency maps [16], and probing tasks [6], aim to improve the interpretability of neural networks. While these methods provide insights into input importance, our work offers a theoretical framework for understanding the internal computations of FFNs, revealing their role as key-value memories [8]. This understanding, in conjunction with empirical analyses like those of Haider et al. [9], can inform more targeted interpretability methods, such as analyzing the folded weights in our derived formulation, and also provides a foundation for more interpretable architectures by explicitly separating knowledge retrieval.